OpenStudy (anonymous):

What is the solution to the system of equations represented by the two equations?

OpenStudy (anonymous):

@nbarrera @william.1233 @inowalst @robtobey

OpenStudy (anonymous):

@marylou004 @JFraser @MakaylaTracy

OpenStudy (anonymous):

Do anyone know

OpenStudy (nbarrera):

Is there any answer choices?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

Solve the following for x:\[\frac{x}{3}=7-2 x \]

OpenStudy (anonymous):

21/3=7

OpenStudy (jfraser):

substitute one value of Y into the other equation for Y and solve

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

Am I Correct @robtobey

OpenStudy (anonymous):

Yes. This system is very slow at this time, so bailing out.

OpenStudy (anonymous):

so what would be the pair

OpenStudy (anonymous):

Would it be ( 1,3)

OpenStudy (anonymous):

Or (3,1)

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

it was (3,1)

OpenStudy (anonymous):

I gotta 100% I did good for myself

OpenStudy (anonymous):

tyurtyuyutyuyutryu

OpenStudy (anonymous):

huh @william.1233

OpenStudy (anonymous):

654564546546545 idk

OpenStudy (anonymous):

oh lol